Suppose you are the manager of a watchmaking firm operating in a competitive market. Your cost of production is given by C = 200 + 2q², where q is the level of output (watches) and C is the total cost.

a) If the price of a watch is $100, how many watches should you produce to maximise profit?

b) Calculate the firm's profits and producer surplus at the equilibrium level of output.

c) What is the difference between profit and producer surplus?

d) If demand increase, what are the steps by which a competitive market ensures increased output?
1
Expert's answer
2014-05-14T11:13:08-0400
TC = 200 + 2q²
a)If the price of a watch is $100, to maximize profit you must produce the quantity for which MR = MC = P.
MC = TC' = 4q
4q = 100
q = 25 watches

b)The firm's profits TP = TR - TC = P*q - TC = 100*25 - 200 - 2*25^2 = 2500 - 200 - 1250 = $1050

Producer surplus is an economic measure of the difference between the amount that a producer of a good receives and the minimum amount that he or she would be willing to accept for the good.
PS = 0.5*100*25 = $1250

c)What is the difference between profit and producer surplus?
TP - PS = 1050 - 1250 = -$200

d)If demand increases, competitive market ensures increased output by the increase of price for good.